OpenStudy (anonymous):

I have an algebra II question that just isn't making sense to me.. I am really unsure about what the question is asking me to do.. so if you could help me work through it , It would be greatly appreciated !! : 13. Determine whether y varies directly with x. If so, find the constant of variation k and write the equation. x | y 4 | 6.4 7 | 11.2 10| 16 13| 20.8

OpenStudy (amtran_bus):

Plot the points and graph.

OpenStudy (anonymous):

Okay i graphed them @AmTran_Bus , now what ??

OpenStudy (amtran_bus):

Let me do it real quick.

OpenStudy (anonymous):

Okay

OpenStudy (amtran_bus):

So I got a straight line, perfectly diagonal.

OpenStudy (amtran_bus):

11.2-6.4=4.8 16-11.2=4.8 20.8-16=4.8

OpenStudy (amtran_bus):

I would say yes